OpenStudy (anonymous):

rewrite the expression in terms of loga(4). loga(64) Dont understand this any help will be greatly appreciated

OpenStudy (campbell_st):

well 64 = 4^3 so you have \[\log_{a}(64) = \log_{a}(4^3)\] now apply the log law for powers to complete the answer

OpenStudy (anonymous):

what's the law of powers???

OpenStudy (campbell_st):

\[\log(x^b) = b \times \log(x)\]

OpenStudy (anonymous):

\[3*\log_a{4} \] is this right

OpenStudy (campbell_st):

that's it

OpenStudy (anonymous):

sweet thanks